# Fluctuations for non-Hermitian dynamics

### v1

@@ Abstract
In the proof, we obtain a quantitative, optimal relaxation at the hard edge, for a broad extension of the Dyson Brownian motion, with a driving noise arbitrarily correlated in space.
@@
University of Arizona
gcipolloni@arizona.edu
@@ 1 Introduction
Here $k$ is an explicit kernel which is smooth off the diagonal, and ||$\cdot$|| ⋅ | is the Euclidean distance in $\mathbb{R}^{2}$.
@@
P. B. was supported by the NSF standard grant DMS-2054851. J. H. was supported by NSF standard grant DMS-2331096 and career grant DMS-2337795, and the Sloan research award.
@@
Here $f$ is a test function supported on $\Omega$, a fixed disk with center 00 and arbitrary radius greater than $1$. Without loss of generality, by polarization, it is enough to consider real test functions $f$, which will be assumed to be in the Sobolev space ${\rm H}_{0}^{2+\varepsilon}(\Omega)$, for a fixed small $\varepsilon>0$, which is defined as the completion of the smooth compactly supported functions $C_{c}^{\infty}(\Omega)$ under the norm
@@
[…]
@@
We have $q_{r}(z){\rm d}z\to\delta_{0}$ as $r\to 0$, so one recovers (1.4). Moreover $q_{r}(z){\rm d}z$ is the pushforward, by stereographic projection, of the uniform probability measure on $\mathscr{S}_{r}$ (the 2-sphere with center 00 and radius $r$). Denoting $(Q_{r}f)(z)=\int f(z-w)q_{r}(w){\rm d}w$ the averaging by this stereographic kernel, Theorem 2.3 means
@@
The mesoscopic covariance $\Gamma_{v}(f,g,\tau)$ from Theorem 2.3 can be directly proved to be positive definite, i.e. $\sum_{1\leqslant i,\,j,\leqslant m}\Gamma_{v}(f_{i},f_{j},|t_{i}-t_{j}|)\geqslant 0$ for any $m\geqslant 1$, functions $f_{i}$ and times $t_{i}$. For $m=2$ this follows easily from Schwarz’s and then Young’s inequality:
@@
We stated Theorem 2.3 only in the bulk regime $|v|<1$, however a similar statement holds when $|v|=1$ if $a\leqslant\varepsilon$, for some very small $N$–independent $\varepsilon>0$. Additionally, for the sake of clarity we stated Theorems 2.22.3 only for two $N$–independent times $s,t$, but the proof also gives an analogous result for multiple, $N$-dependent, times.
@@
[…]
@@
To handle microscopic scales, [CipErdSch2023] proved that the smallest singular values of Hermitized matrices corresponding to distant spectral points are independent, by technically difficult variants of the dynamical method [LanSosYau2019]. Theorem 4.1 proves that this independence holds in much greater generality, under the assumptions of some weak local law instead of rigidity of the particles, and natural bounds on eigenvector overlaps, which are model-dependent and follow in our case from important estimates in [CipErdSch2023]. This theorem could be applied to a wide range of non-Hermitian models with non-trivial mean and a variance profile, or directed graphs. Furthermore, this method has potential to apply to many problems involving eigenvalues statistics close to an hard edge.
@@
    It initiates the study of a natural extension of the Dyson Brownian motion, as it covers relaxation for arbitrary, correlated martingales driving individual particles, see (4.5) and the minimal assumption (4.7).
@@
    It provides the optimal, submicroscopic error estimates (up to subpolynomial terms). In particular ($s_{1}$ and $s_{1}^{\prime}$ are the particle closest to the hard edge)
@@
    for any $\bm{s}(0),\bm{s}^{\prime}(0)$ satisfying a weak local law and $\bm{s}(t),\bm{s}^{\prime}(t)$ following the extended Dyson Brownian Motion dynamics.
A key tool for the above relaxation estimate is an observable introduced in [Bou2022]. However the proof of (2) is considerably simpler than the optimal relaxation of the eigenvalue gaps in [Bou2022], as it bypasses a decomposition into short and long range interactions that has been customary in the dynamics approach [ErdYau2015, BouYau2017, LanSosYau2019], and does not require any maximum principle.
@@
A surprising aspect of Theorems 2.22.3 is the remarkably simple limiting space-time logarithmic correlation, despite its emergence from the superposition of the complicated covariances from a family of resolvents (from the Hermitization). This simplicity can be interpreted, in Corollary 2.11, as a statement on space-time correlations between eigenvector overlaps. Indeed Equation (2.9) follows from the combination of Theorem 2.3 and the non-Hermitian dynamics (1.5), at equilibrium. Out of equilibrium, such a quartic decay of correlations seems out of reach: To exhibit overlaps at two distinct times, we need to consider not only forward but also backward dynamics, which are tractable only at equilibrium (see Subsection 5.3).
@@
For two quantities $X$ and $Y$ depending on $N$, we write that $X=\operatorname{O}(Y)$ or $X\lesssim Y$ if there exists some universal constant $C>0$ such that $|X|\leqslant CY$ We write $X=\operatorname{o}(Y)$, or $X\ll Y$ if the ratio $|X|/Y\rightarrow\infty$ as $N$ goes to infinity. We write $X\asymp Y$ if there exists a universal constant $C>0$ such that $Y/C\leqslant|X|\leqslant CY$. We denote $\llbracket a,b\rrbracket=[a,b]\cap{\mathbb{Z}}$ and $\llbracket n\rrbracket=\llbracket 1,n\rrbracket$. Furthermore, for a matrix $\mathbb{C}^{d\times d}$ we denote its normalized trace by $\langle A\rangle:=d^{-1}\mathrm{Tr}[A]$. Because of the Hermitization, in this paper $d=2N$.
@@
with $\eta_{0}=N^{-1-\delta_{0}}$, $\eta_{c}=N^{-1+\delta_{1}},$ for some $\delta_{0},\delta_{1}>0$. In order to keep the notation for the proof of the macroscopic and mesoscopic CLT unified, with a slight abuse of notation, here we use the convention that $f_{v,a}=f$ when $a=0$.
The integral $J_{T}$ can be easily seen to be negligible (see e.g. the proof of [CipErdSch2021bis, Lemma 3] and [AEK21, Theorem 2.3]). Next, we show that the very small $\eta$–regime $I_{0}^{\eta_{0}}$ is negligible using smoothing inequalities for the smallest singular value of $X-z$ (see Section 3.2). The regime $I_{\eta_{0}}^{\eta_{c}}$ will also be negligible (see Section 3.3), for this term we will rely on the asymptotical independence of the small singular values of $X_{t}-z_{1}$, $X_{t}-z_{2}$ for $z_{1},z_{2}$ sufficiently away from each other. Finally, we will compute high moments of $I_{\eta_{c}}^{T}$ showing that this is the regime where the order one contribution to the lhs. of (3.5) comes from (see Section 3.4).
@@
By the lower tail estimate for the smallest singular value of $X-z$ from [TaoVu2010, Theorem 3.2] (see also [CipErdSch2020, Equation (4a)]), we readily conclude that the contribution of this regime is negligible. In particular, we have (cf. [CipErdSch2023, Lemma 4.4])
@@
For fixed $c,\omega_{p}>0$, let $\delta_{0},\delta_{1}$ be small enough constants. Then uniformly in $\lvert z_{l}\rvert\leqslant 1-c$, $\lvert z_{1}-z_{2}\rvert\geqslant N^{-1/2+\omega_{p}}$, and $\eta_{1},\eta_{2}\in[N^{-1-\delta_{0}},N^{-1+\delta_{1}}]$, we have [^2] We state this result for only two different $z$’s for simplicity, but a similar result holds for any finite product of resolvents such that $\min_{i\neq j}|z_{i}-z_{j}|\gg N^{-1/2}$.
$\mathbb{E}\langle G_{t}^{z_{1}}(\mathrm{i}\eta_{1})\rangle\langle G_{t}^{z_{2}}(\mathrm{i}\eta_{2})\rangle=\mathbb{E}\langle G_{t}^{z_{1}}(\mathrm{i}\eta_{1})\rangle\mathbb{E}\langle G_{t}^{z_{2}}(\mathrm{i}\eta_{2})\rangle+\operatorname{O}\left(N^{-{\varepsilon}}\right).$ (3.8)
@@
This is straightforward by expanding the square in $\mathbb{E}|I_{\eta_{0}}^{\eta_{c}}(f_{v,a},t)|^{2}$, using (3.8) and $\int|\Delta f|\lesssim 1$ (see the proof of [CipErdSch2023, Lemma 4.4]). ∎
@@ 4 Independence of singular values
$\nu,\omega\ll\omega_{t}\ll\omega_{K}\ll\tilde{\omega}.$
@@
  • $\nu$. The local law holds on scale $\varphi/N$ (Equation (4.2)).
@@
  • $\tilde{\omega}$. These eigenvector products are of size at most $N^{-\tilde{\omega}}$ (Equation (4.3)).
@@
Let $X$ be a $N\times N$ matrix with complex entries, and consider the singular values $\lambda_{i}^{z}$ of $X-z$, for $z\in I:=\{z_{1},\dots,z_{q}\}$, $q$ fixed, $z_{1},\dots,z_{q}\in\mathbb{C}$ possibly $N$-dependent.
Assume the following local law on scale $\varphi/N$ holds. There is a $c>0$ and $N_{0}$ such that for any $N\geqslant N_{0}$ and $z\in I$ there exists $s=s^{z}$, the Stieltjes transform of a deterministic probability measure $\sigma=\sigma^{z}$, such that $c\leqslant|{\rm Im}\,s(w)|\leqslant c^{-1}$, $|s(w)|+|s^{\prime}(w)|\leqslant c^{-1}$ for all $|w|\leqslant c$, and such that for any ${\rm Im}w\geqslant\varphi/N$, $|w|\leqslant c$ we have
$\left|\sum_{i=1}^{N}\frac{1}{\lambda_{i}^{z}-w}-s^{z}(w)\right|\leqslant\frac{1}{\sqrt{N{\rm Im}w}}.$ (4.2)
Let $X_{t}$ be the solution of (1.5), and for $z\in I$ denote the singular values of $X_{t}-z$ by $\lambda_{i}^{z}(t)$. Let ${\bm{u}}_{i}^{z}(t),{\bm{v}}_{i}^{z}(t)$ be the left and right singular vectors of $X_{t}-z$, respectively, and assume that there exists $\omega_{K}\ll\tilde{\omega}$ such that for any $z,w$ distinct in $I$ and $D>0$ we have
[…]
for a fixed $T>0$. Let $(X^{(p)})_{1\leqslant p\leqslant q}$ be independent Ginibre matrices evolving along (1.5), and denote the singular values of $X^{(p)}$ by $\mu_{i}^{(p)}(t)$. Then there exists a coupling and constants $\omega\ll\omega_{t}\ll\omega_{K}$ so that for any $z_{p}\in I$ we have
$\mathbb{P}\left(\big{|}\rho_{t}(0)\lambda_{i}^{z_{p}}(t)-\rho_{\mathrm{sc}}^{(t)}(0)\mu_{i}^{(p)}(t)\big{|}\leqslant N^{-1-\omega},\ \forall 1\leqslant i\leqslant N^{\omega}\ {\rm and}\ t\in[N^{-1+\omega_{t}},T]\right)\geqslant 1-N^{-D}.$ (4.4)
Here $\rho_{t}$ (resp. $\rho_{\mathrm{sc}}^{(t)}$) is the density of $\sigma^{z_{p}}\boxplus\mu_{\rm sc}^{(t)}$ (resp. $\mu_{\mathrm{sc}}^{(1)}\boxplus\mu_{\rm sc}^{(t)}$) (see the discussion in Section 4.4 for these notations for measures and their free convolutions).
@@
By a standard application of a Green’s function comparison (GFT) argument, it is enough to prove (3.8) for matrices with an additional Gaussian component of size $\sqrt{t}$, with $t\leqslant N^{-1/2-\varepsilon}$, for any small fixed $\varepsilon>0$ (see e.g. [CipErdSch2023, Lemma 7.5]). We will thus prove (3.8) for matrices having a Gaussian component of size $\sqrt{T}$, with $T=N^{-1+\omega_{t}}$, for some fixed small $\omega_{t}>0$.
For $|z_{1}-z_{2}|\geqslant N^{-1/2+\omega_{p}}$, by [CipErdSch2022bis, Theorem 3.1], there exist $\tilde{\omega},\omega_{K}$ such that (4.3) is satisfied for a fixed time. To show the bound (4.3) uniformly in time it is enough to use a standard grid argument as a consequence of the fact that the bound on eigenvectors is inherited by a bound on product of resolvents (see e.g. the proof of [CipErdSch2023, Lemma 7.9]) which is Hölder in time. Additionally, the single resolvent local law [CipErdSch2021, Theorem 3.1] implies (4.2). This shows that the assumptions of Theorem 4.1 are satisfied, and (4.4) holds. Given (4.4) as an input (cf. [CipErdSch2023, Lemma 7.6–7.7]), using rigidity (cf. [CipErdSch2023, Proposition 7.3]), the proof of (3.8) is completely analogous to the proof of [CipErdSch2023, Proposition 7.2] and so omitted. ∎
@@
In this section, the continuous martingales $b_{i}^{s}(t)$ and $b_{i}^{r}(t)$ are realized on a common probability space with a common filtration $\mathcal{F}_{t}$. They do not need to be Brownian motions. They always satisfy the symmetry $b^{s}_{-i}(t)=-b^{s}_{i}(t)$$b^{r}_{-i}(t)=-b^{r}_{i}(t)$, and the quadratic variation bound

### v2

# arXiv:2409.02902v2[math.PR] 02 Mar 2026

@@ Abstract
In the proof, we obtain a quantitative relaxation at the hard edge, for an extension of the Dyson Brownian motion, with a driving noise arbitrarily correlated in space.
[…]
@@
University of Rome Tor Vergata
cipolloni@axp.mat.uniroma2.it
@@ 1 Introduction
Here $k$ is an explicit kernel which is smooth off the diagonal, and $|\cdot|$ is the Euclidean distance in $\mathbb{R}^{2}$.
@@
P. B. was supported by the NSF standard grant DMS-2054851. J. H. was supported by NSF standard grant DMS-2331096 and career grant DMS-2337795, and the Sloan research award. G. C. is partially supported by the MUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C18000100006. We also thank the referees for their excellent review work.
@@
Here $f$ is a test function supported on $\Omega$, a fixed disk with center 0 and arbitrary radius greater than $1$. Without loss of generality, by polarization, it is enough to consider real test functions $f$, which will be assumed to be in the Sobolev space ${\rm H}_{0}^{2+\varepsilon}(\Omega)$, for a fixed small $\varepsilon>0$, which is defined as the completion of the smooth compactly supported functions $C_{c}^{\infty}(\Omega)$ under the norm
@@
Let $X_{t}$ be the solution of (1.5), with $X$ satisfying Assumption 2.1. Consider real valued $f,g\in H_{0}^{2+\varepsilon}(\Omega)$. Then for any fixed $s\leqslant t$, $L_{N}(f,t),L_{N}(g,s)$ converge jointly in distribution to centered Gaussian random variables $(L(f,t)$, $L(g,s))$ with covariance
@@
We have $q_{r}(z){\rm d}z\to\delta_{0}$ as $r\to 0$, so one recovers (1.4). Moreover $q_{r}(z){\rm d}z$ is the pushforward, by stereographic projection, of the uniform probability measure on $\mathscr{S}_{r}$ (the 2-sphere with center 0 and radius $r$). Denoting $(Q_{r}f)(z)=\int f(z-w)q_{r}(w){\rm d}w$ the averaging by this stereographic kernel, Theorem 2.3 means
@@
The mesoscopic covariance $\Gamma_{v}(f,g,\tau)$ from Theorem 2.3 can be directly proved to be positive definite, i.e. $\sum_{1\leqslant i,\,j\leqslant m}\Gamma_{v}(f_{i},f_{j},|t_{i}-t_{j}|)\geqslant 0$ for any $m\geqslant 1$, functions $f_{i}$ and times $t_{i}$. For $m=2$ this follows easily from Schwarz’s and then Young’s inequality:
@@
We stated Theorem 2.3 only in the bulk regime $|v|<1$, however a similar statement holds when $|v|=1$ if $a\leqslant\varepsilon$, for some very small NN–independent $\varepsilon>0$. Additionally, for the sake of clarity we stated Theorems 2.22.3 only for two NN–independent times $s,t$, but the proof also gives an analogous result for multiple, $N$-dependent, times.
@@
$\begin{split}&\phantom{{}={}}\int_{[s_{1},s_{2}]\times[t_{1},t_{2}]}\mathbb{E}\Big[\sum_{i,j}f_{v,a}(\sigma_{i}(s))g_{v,a}(\sigma_{j}(t))\Big(\frac{\mathscr{O}_{ii}(s)}{N}-c_{v}\Big)\Big(\frac{\mathscr{O}_{jj}(t)}{N}-c_{v}\Big)\Big]{\rm d}s{\rm d}t\\ &=\int_{[s_{1},s_{2}]\times[t_{1},t_{2}]}\int_{\mathbb{C}^{2}}f_{v,a}(z)g_{v,a}(w)\frac{c_{v}^{2}}{(c_{v}|t-s|+|z-w|^{2})^{2}}\frac{{\rm d}z{\rm d}w}{\pi^{2}}{\rm d}s{\rm d}t+\operatorname{O}(N^{-4a-{\varepsilon}}).\end{split}$ (2.9)
@@
To handle microscopic scales, [CipErdSch2023] proved that the smallest singular values of Hermitized matrices corresponding to distant spectral points are independent, by technically difficult variants of the dynamical method [LanSosYau2019]. This independence result from [CipErdSch2023] would be sufficient for that part of the proof of our main theorems, but we present a self-contained general Theorem 4.1, which gives this independence in much greater generality, under the assumptions of some weak local law instead of rigidity of the particles, and natural bounds on eigenvector overlaps, which are model-dependent and follow in our case from important estimates in [CipErdSch2023]. This theorem could be applied to a wide range of non-Hermitian models with non-trivial mean and a variance profile, or directed graphs. Furthermore, this method has potential to apply to many problems involving eigenvalues statistics close to a hard edge.
@@
    It initiates the study of a natural extension of the Dyson Brownian motion, as it covers relaxation for arbitrary, correlated martingales driving individual particles, see (4.5) and the minimal assumption (4.7). Although, currently, this new correlated setting does not clearly connect to random matrices, dynamics with general colored noise is of general interest in stochastic (partial) differential equations, see for example Section 7.2.4 in the classical book [DaPZab] for general theory, and [DaPGat] for the stochastic Burgers equation. Proposition 4.6 provides an example of universal local convergence (i.e. independent on the initial condition), for singular SDE coefficients and essentially arbitrary noise structure in space (but white in time).
@@
    We obtain the following submicroscopic error estimates (up to subpolynomial terms in $N$) for the relaxation under the sole assumptions of an initial local law and arbitrary noise. In particular ($s_{1}$ and $s_{1}^{\prime}$ are the particles closest to the hard edge)
@@
    for any $\bm{s}(0),\bm{s}^{\prime}(0)$ satisfying a weak local law with respect to the same deterministic density, and $\bm{s}(t),\bm{s}^{\prime}(t)$ following the extended (correlated) Dyson Brownian Motion dynamics. We refer to (4.61) for an explanation of this error term in this very general setting. When the underlying Brownian motions are independent and the initial conditions satisfy a stronger initial rigidity assumption, relaxation holds with a smaller error term as explained in Remark 4.19.
A key tool for the above relaxation estimate is an observable introduced in [Bou2022], defined in (4.66). However the proof of (2) is considerably simpler than the optimal relaxation of the eigenvalue gaps in the bulk of the spectrum [Bou2022], as it bypasses a decomposition into short and long range interactions that has been customary in the dynamics approach [ErdYau2015, BouYau2017, LanSosYau2019], and does not require any maximum principle. Our proof of the hard edge relaxation is therefore closer in spirit to the proof of soft edge relaxation given in [Bou2022]: It demonstrates that the sole analysis of a key averaged observable gives microscopic statistics at the hard edge, by fully exploiting the symmetry of the particles. After [Bou2022] the possibility of such a proof at the hard edge was unclear, because the interparticle distance $N^{-1}$ at the hard edge and in the bulk coincide.
Moreover, in the considered setting of arbitrary correlations for the noise, the observable from [Bou2022] is still the key to quantitative relaxation, although meaningful changes are needed compared to the case of independent noises: in [Bou2022], rigidity (a local law with error $(N\eta)^{-1}$) is preserved through the dynamics, which eventually yields the small error term $\frac{1}{N^{2}t}$ for the relaxation of gaps in the bulk of the spectrum, while in our setting only a weaker local law (with error $(N\eta)^{-1/2}$, see Proposition 4.15) is preserved, which is responsible for the weaker bound $\frac{1}{N\cdot\sqrt{Nt}}$ at the hard edge. This deteriorated bound would also be present if arbitrary martingales were considered in [Bou2022], due, for example, to a larger bracket term in the proof of [Bou2022, Lemma 3.10].
@@
A surprising aspect of Theorems 2.22.3 is the remarkably simple limiting space-time logarithmic correlation, despite its emergence from the superposition of the complicated covariances from a family of resolvents (from the Hermitization). This simplicity can be interpreted, in Corollary 2.11, as a statement on space-time correlations between eigenvector overlaps. Indeed, Equation (2.9) follows from the combination of Theorem 2.3 and the non-Hermitian dynamics (1.5), at equilibrium. Out of equilibrium, such a quartic decay of correlations seems out of reach: To exhibit overlaps at two distinct times, we need to consider not only forward but also backward dynamics, which are tractable only at equilibrium (see Subsection 5.3).
@@
For two quantities $X$ and $Y$ depending on $N$, we write that $X=\operatorname{O}(Y)$ or $X\lesssim Y$ if there exists some universal constant $C>0$ such that $|X|\leqslant CY$ . We write $X=\operatorname{o}(Y)$, or $X\ll Y$ if the ratio $|X|/Y\rightarrow\infty$ as $N$ goes to infinity. We write $X\asymp Y$ if there exists a universal constant $C>0$ such that $Y/C\leqslant|X|\leqslant CY$. We denote $\llbracket a,b\rrbracket=[a,b]\cap{\mathbb{Z}}$ and $\llbracket n\rrbracket=\llbracket 1,n\rrbracket$. Furthermore, for a matrix $\mathbb{C}^{d\times d}$ we denote its normalized trace by $\langle A\rangle:=d^{-1}\mathrm{Tr}[A]$. Because of the Hermitization, in this paper $d=2N$.
@@
with $\eta_{0}=N^{-1-\delta_{0}}$, $\eta_{c}=N^{-1+\delta_{1}},$ for some $\delta_{0},\delta_{1}>0$, and $T=N^{100}$. In order to keep the notation for the proof of the macroscopic and mesoscopic CLT unified, with a slight abuse of notation, here we use the convention that $f_{v,a}=f$ when $a=0$.
The integral $J_{T}$ can be easily seen to be negligible (see e.g. the proof of [CipErdSch2021bis, Lemma 3] and [AEK21, Theorem 2.3]). Next, we show that the very small $\eta$–regime $I_{0}^{\eta_{0}}$ is negligible using smoothing inequalities for the smallest singular value of XzX-z (see Section 3.2). The regime $I_{\eta_{0}}^{\eta_{c}}$ will also be negligible (see Section 3.3), for this term we will rely on the asymptotical independence of the small singular values of $X_{t}-z_{1}$, $X_{t}-z_{2}$ for $z_{1},z_{2}$ sufficiently away from each other. Finally, we will compute high moments of $I_{\eta_{c}}^{T}$ showing that this is the regime where the order one contribution to the lhs. of (3.5) comes from (see Section 3.4).
@@
By the lower tail estimate for the smallest singular value of XzX-z from [TaoVu2010, Theorem 3.2] (see also [CipErdSch2020, Equation (4a)]), we readily conclude that the contribution of this regime is negligible. In particular, we have (cf. [CipErdSch2023, Lemma 4.4])
@@
Fix any small $c,\omega_{p}>0$, let $\delta_{0},\delta_{1}$ be small enough constants. Then uniformly in $\lvert z_{l}\rvert\leqslant 1-N^{-c}$, $\lvert z_{1}-z_{2}\rvert\geqslant N^{-1/2+\omega_{p}}$, and $\eta_{1},\eta_{2}\in[N^{-1-\delta_{0}},N^{-1+\delta_{1}}]$, we have [^2] We state this result for only two different $z$’s for simplicity, but a similar result holds for any finite product of resolvents such that $\min_{i\neq j}|z_{i}-z_{j}|\gg N^{-1/2}$.
$\mathbb{E}\langle G_{t}^{z_{1}}(\mathrm{i}\eta_{1})\rangle\langle G_{t}^{z_{2}}(\mathrm{i}\eta_{2})\rangle=\mathbb{E}\langle G_{t}^{z_{1}}(\mathrm{i}\eta_{1})\rangle\mathbb{E}\langle G_{t}^{z_{2}}(\mathrm{i}\eta_{2})\rangle+\operatorname{O}\left(N^{-{\varepsilon}}\right),$ (3.8)
for some ${\varepsilon}>0$.
@@
Using that the regime $|z_{1}-z_{2}|<N^{-1/2+\omega_{p}}$ is negligible (see e.g. [CipErdSch2023, Eq. (4.12)] and [CipErdSch2022bis, Eq. (4.8)]), this is straightforward by expanding the square in $\mathbb{E}|I_{\eta_{0}}^{\eta_{c}}(f_{v,a},t)|^{2}$, using (3.8) and $\int|\Delta f|\lesssim 1$ (see the proof of [CipErdSch2023, Lemma 4.4]). We point out that in the proof of [CipErdSch2023, Lemma 4.4] it is shown that the regime $|z_{l}|\geqslant 1-N^{-c}$ is negligible (see the paragraph below [CipErdSch2023, Eq. (4.11)]) and so that to conclude (3.7) it is indeed enough to use (3.8) for $|z_{l}|\leqslant 1-N^{-c}$. ∎
@@ 4 Independence of singular values
$\nu,\omega\ll\omega_{t}\ll\omega_{K}\ll\widetilde{\omega}$.
@@
  • $\nu$. For $\varphi$ from (4.1), the local law holds on scale $\varphi/N$ (see Equation (4.2)).
@@
  • $\widetilde{\omega}$. These eigenvector products are of size at most $N^{-\widetilde{\omega}}$ (Equation (4.3)).
@@
Let $X$ be a $N\times N$ matrix with complex entries, and consider the singular values $\lambda_{i}^{z}$ of XzX-z, for $z\in I:=\{z_{1},\dots,z_{q}\}$, $q$ fixed, $z_{1},\dots,z_{q}\in\mathbb{C}$ possibly $N$-dependent.
Assume the following local law on scale $\varphi/N$ holds. There is a $c>0$ and $N_{0}$ such that for any $N\geqslant N_{0}$ and $z\in I$ there exists $m=m^{z}$, the Stieltjes transform of a deterministic probability measure $\zeta=\zeta^{z}=\rho^{z}_{0}(x){\rm d}x$, such that $c\leqslant|{\rm Im}\,m(w)|\leqslant c^{-1}$, $|m(w)|+|\partial_{w}m(w)|\leqslant c^{-1}$ for all $|w|\leqslant c$, and such that for any ${\rm Im}w\geqslant\varphi/N$, $|w|\leqslant c$ we have
$\left|\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\lambda_{i}^{z}-w}-m^{z}(w)\right|\leqslant\frac{1}{\sqrt{N{\rm Im}w}}.$ (4.2)
Let $X_{t}:=X+N^{-1/2}B_{t}$, with $(B_{t})_{ij}$ being i.i.d. standard complex Brownian motions, and for $z\in I$ denote the singular values of $X_{t}-z$ by $\lambda_{i}^{z}(t)$. Let ${\bm{u}}_{i}^{z}(t),{\bm{v}}_{i}^{z}(t)$ be the left and right singular vectors of $X_{t}-z$, respectively, and assume that there exists $\omega_{K}\ll\widetilde{\omega}$ such that for any $z,w$ distinct in $I$ and $D>0$ we have
$\mathbb{P}\Big(\big|\langle{\bm{u}}_{i}^{z}(t),{\bm{u}}_{j}^{w}(t)\rangle\big|+\big|\langle{\bm{v}}_{i}^{z}(t),{\bm{v}}_{j}^{w}(t)\rangle\big|\leqslant N^{-\widetilde{\omega}},\quad\forall\,1\leqslant i,j\leqslant N^{\omega_{K}}\,\,\mathrm{and}\,\,t\in[0,\mathcal{T}]\Big)\geqslant 1-N^{-D},$ (4.3)
for a fixed $\mathcal{T}>0$. Let $(X^{(p)})_{1\leqslant p\leqslant q}$ be independent Ginibre matrices, and denote the singular values of $X^{(p)}$ by $\mu_{i}^{(p)}$. Then for any time $t\in[N^{-1+\omega_{t}},\mathcal{T}]$ there exists a coupling and constants $\omega\ll\omega_{t}\ll\omega_{K}$ so that for any $z_{p}\in I$ we have
$\mathbb{P}\left(\big|\rho_{t}(0)\lambda_{i}^{z_{p}}(t)-\rho_{\mathrm{sc}}(0)\mu_{i}^{(p)}\big|\leqslant N^{-1-\omega},\ \forall 1\leqslant i\leqslant N^{\omega}\right)\geqslant 1-N^{-D}.$ (4.4)
Here $\rho_{t}$ is the density of $\zeta^{z_{p}}\boxplus\mu_{\rm sc}^{(t)}$ (see the discussion in Section 4.4 for these notations for measures and their free convolutions), and $\rho_{\rm sc}$ is the density of the semi-circle law.
@@
By a standard application of a Green’s function comparison (GFT) argument, it is enough to prove (3.8) for matrices with an additional Gaussian component of size $\sqrt{t}$, with $t\leqslant N^{-1/2-\varepsilon}$, for a small fixed $\varepsilon>0$ (see e.g. [CipErdSch2023, Lemma 7.5]). We will thus prove (3.8) for matrices having a Gaussian component of size $\sqrt{\mathcal{T}}$, with $\mathcal{T}=N^{-1+\omega_{t}}$, for some fixed small $\omega_{t}>0$. We now show that for matrices with a Gaussian component of size $\sqrt{\mathcal{T}}$ the desired result follows from Theorem 4.1. We point out that Theorem 4.1 is stated under the assumption $c\leqslant|{\rm Im}\,m(w)|\leqslant c^{-1}$, $|m(w)|+|\partial_{w}m(w)|\leqslant c^{-1}$, for an $N$-independent $c>0$. When $1-|z|\sim N^{-o(1)}$, this assumption is not satisfied as $|{\rm Im}\,m(w)|\sim N^{-o(1)}$ and $|\partial_{w}m(w)|\leqslant N^{o(1)}$ (see e.g. [CipErdXu2025, Eq. (3.10)]). However, the conclusion in (4.4) holds also under this slightly weaker assumptions as stated in Remark 4.4 below.
For $|z_{1}-z_{2}|\geqslant N^{-1/2+\omega_{p}}$, by [CipErdSch2022bis, Theorem 3.1] and [CipErdSch2023, Lemma 7.9] (see also [CipErdXu2025, Corollary 3.6]), there exist $\widetilde{\omega},\omega_{K}$ such that (4.3) is satisfied for a fixed time. To show the bound (4.3) uniformly in time it is enough to use a standard grid argument as a consequence of the fact that the bound on eigenvectors is inherited by a bound on product of resolvents (see e.g. the proof of [CipErdSch2023, Lemma 7.9]) which is Hölder in time. Additionally, the single resolvent local law [CipErdSch2021, Theorem 3.1] implies (4.2). This shows that the assumptions of Theorem 4.1 are satisfied, and (4.4) holds. Given (4.4) as an input (cf. [CipErdSch2023, Lemma 7.6–7.7]), using rigidity (cf. [CipErdSch2023, Proposition 7.3]), the proof of (3.8) is completely analogous to the proof of [CipErdSch2023, Proposition 7.2] and so omitted. ∎
@@
In this section, the continuous martingales $b_{i}^{s}(t)$ and $b_{i}^{r}(t)$ are realized on a common probability space with a common filtration $\mathcal{F}_{t}$. They do not need to be Brownian motions. They always satisfy the symmetry $b^{s}_{-i}(t)=-b^{s}_{i}(t)$, $b^{r}_{-i}(t)=-b^{r}_{i}(t)$, and the quadratic variation bound